# GEM: Gaussian Embedding Modeling for Out-of-Distribution Detection in GUI Agents

### v2

@@ Abstract
Graphical user interface (GUI) agents have recently emerged as an intriguing paradigm for human-computer interaction, capable of automatically executing user instructions to operate intelligent terminal devices. However, when encountering out-of-distribution (OOD) instructions that violate environmental constraints or exceed the current capabilities of agents, GUI agents may suffer task breakdowns or even pose security threats. Therefore, effective OOD detection for GUI agents is essential. Traditional OOD detection methods perform suboptimally in this domain due to the complex embedding space and evolving GUI environments. In this work, we observe that the in-distribution input semantic space of GUI agents exhibits a clustering pattern with respect to the distance from the centroid. Based on the finding, we propose GEM, a novel method based on fitting a Gaussian mixture model over input embedding distances extracted from the GUI Agent that reflect its capability boundary. Evaluated on eight datasets spanning smartphones, computers, and web browsers, our method achieves an average accuracy improvement of 23.70% over the best-performing baseline. Analysis verifies the generalization ability of our method through experiments on nine different backbones. The codes are available at https://github.com/Wuzheng02/GEM-OODforGUIagents.
1 Introduction
Recently, graphical user interface (GUI) agents (nguyen2024gui, ; wang2024gui, ; zhang2024large, ; liu2025llm, ; shi2025towards, ) have emerged as an intriguing paradigm to human-computer interaction, capable of autonomously executing user instructions and performing human-like control on intelligent terminal devices such as smartphones, computers, and web browsers. The common approach to building GUI agents involves post-training multimodal large language models (MLLMs) using high-quality trajectory data to enhance key task capabilities such as perception li2023blip ; wang2023image , reasoning yao2022react ; wei2022chain ; zhang2023multimodal , and reflection (liu2025infiguiagent, ; hu2025guiding, ).
@@ 1 Introduction
Despite notable advances in instruction following, GUI agents remain vulnerable to out-of-distribution (OOD) risks—executing instructions that violate environmental constraints (e.g., non-existent functions or applications) or exceed the agent’s current capabilities. These failures can lead to task breakdowns or even pose security threats. In real-world applications, OOD risks for GUI agents come with two primary forms: (i) Internalization-OOD, where the agent operates in domain-specific environments (e.g., a particular type of smartphones or vehicle cabins) but incorrectly assumes the presence of unsupported capabilities; (ii) Extrapolation-OOD, where the agent encounters instructions tied to dynamic or evolving environments (e.g., new third-party applications).
@@
Popular OOD detection methods can be broadly categorized malinin2018predictive ; yang2024generalized into two types: embedding-based approaches lee2018simple ; sun2022out and model uncertainty-based approaches hendrycks2016baseline ; gal2016dropout . These methods have demonstrated effectiveness in traditional (M)LLM tasks such as summarization ren2022out , visual question answering (VQA) kervadec2021roses , mathematical reasoning (wang2024embedding, ), and text generation wu2022multi . However, as we will show in Section 3, directly applying these existing OOD detection techniques to the GUI agent domain results in suboptimal performance. Notably, even with the classification threshold calibrated to correctly identify 95% of in-distribution (ID) tasks, the best-performing method still misclassifies between 31.75% and 48.92% of OOD tasks as ID.
@@
$\bullet$ Complex Embedding Space: The inputs to GUI agents are inherently complex (ma2024comprehensive, ). Concretely, GUI screens typically contain densely populated UI components zhang2024large , leading to much higher information density than traditional MLLM tasks. Besides, the diversity of user instructions further complicates the inference of user intent. This complexity substantially increases the difficulty of effective OOD detection for GUI agents.
$\bullet$ Evolving Environments: GUI environments frequently change due to system upgrades or the installation of new third-party applications. As a result, GUI agents must continually adapt their capabilities (jiang2025appagentx, ; wang2025mobileE, ; liu2025learnact, ). This evolving nature complicates OOD detection by demanding both accurate capability assessment and temporal adaptability. Consequently, reliance on static external classifiers becomes increasingly impractical.
To address the challenges above, we propose GEM, a novel method based on fitting a Gaussian mixture model (GMM) reynolds2009gaussian over input embedding distances extracted from the GUI Agent that reflect its capability boundary. Specifically, we first extract the input embeddings from the encoder layer of the GUI agent on its training data, and fit them into a high-dimensional hypersphere. We then compute the L2-norm distances of all embeddings relative to the centroid of the hypersphere. Under bayesian information criterion (BIC) neath2012bayesian supervision, we fit a GMM to these distances and define the OOD detection boundary as a configurable number of standard deviations away from each GMM cluster center. Experiments across eight GUI agent datasets spanning smartphones, computers, and web browsers platforms show that our method consistently outperforms all traditional OOD detection methods.
@@
(ii) We propose a novel GMM-based approach that leverages input embedding distances from MLLMs for OOD detection. Our method achieves an average accuracy improvement of 23.70% over the best performing baseline across eight datasets spanning three platform types, consistently outperforming all baseline methods.
(iii) We identify the unique characteristics of OOD detection for GUI agents and provide a comprehensive comparison between our approach and traditional methods, offering new insights into this emerging research area.
2 Related Work
GUI Agents
With the rapid advancement of MLLMs (liu2023visual, ; achiam2023gpt, ; bai2025qwen2, ), numerous cross-platform GUI agent foundation models (hong2024cogagent, ; wu2024atlas, ; qin2025ui, ; xu2024aguvis, ) have emerged, capable of automatically executing user instructions across smartphones, computers, and web browsers. For the technical framework, prompt-based approaches li2024appagent ; wang2025mobileE ; jiang2025appagentx have been developed, leveraging closed-source models to construct agent systems that fulfill user instructions. Additionally, researchers have also explored pre-training ma2024comprehensive ; wu2025smoothing , supervised fine-tuning (SFT) zhang2023you ; wu2025reachagent ; qin2024tool , and reinforcement learning (RL) zhoudigirl ; liu2024autoglm ; lai2024autowebglm ; wang2024distrl ; lu2025ui ; xia2025gui techniques to further enhance the ability of GUI agents to autonomously complete user instructions. Regardless of their advancements, GUI agents inevitably face capability limitations. In complex and dynamic real-world environments chengkairos ; chen2025obvious ; lee2025safeguarding , GUI agents are prone to encountering OOD situations. As a result, robust OOD detection mechanisms for GUI agents remain a critical and valuable area of research.
OOD Detection in MLLMs
For OOD detection in MLLMs, it is necessary to jointly consider both visual and textual modalities, in contrast to traditional computer vision OOD detection, which only requires modeling the visual modality liu2020energy ; ren2019likelihood , or LLMs OOD detection, which focuses on textual modality wang2024embedding ; ren2022out . In the context of MLLMs, some researchers have proposed using maximum concept matching ming2022delving to characterize OOD uncertainty, while others have approached the problem by training models on ID datasets and classifying whether an input image belongs to an unknown category wang2023clipn . However, OOD detection for MLLMs remains a challenging research area dong2024multiood ; lu2024recent . Furthermore, for MLLM-based GUI agents, task scenarios are significantly more complex and dynamic shi2025towards compared to traditional MLLM tasks, making OOD detection even more difficult.
3 Investigating the Challenge of OOD Dection for GUI Agents
In this section, we conduct pilot experiments to evaluate the effectiveness of popular OOD detection methods for GUI agents and understand the challenges. We will first provide the problem formulation, and then present the experimental settings followed by the key results and analysis of the pilot study.
3.1 Problem Formulation for OOD Detection for GUI Agents
A GUI agent $\mathcal{F}$ is initially trained on an ID dataset $\mathcal{D}_{\text{ID}}=\{(s_{i},x_{i})\}_{i=1}^{k}$, consisting of $k$ pairs of screenshots $s_{i}$ and user instructions $x_{i}$. The GUI agent $\mathcal{F}$ learns operational knowledge of operating systems through methods such as SFT and RL on $\mathcal{D}_{\text{ID}}$.
After deployment on the device of a user, the GUI agent $\mathcal{F}$ receives a instruction $x$ and captures the current device screenshot $s_{t}$. $\mathcal{F}$ then constructs a prompt that combines the screenshot $s_{t}$ and the instruction $x$, which is subsequently used to predict an action $a_{t}$. Formally, at each time step $t$ ($0\leq t\leq T$), the process can be expressed as: $a_{t}=\mathcal{F}(s_{t},x)$ where $s_{t}$ is the screenshot at time step $t$, $x$ is the instruction, and $a_{t}$ is the action predicted by the agent at time step $t$.
Then this action $a_{t}$ is executed, leading to a new screenshot $s_{t+1}$. $\mathcal{F}$ evaluates whether the instruction $x$ has been completed. If $x$ is not completed, the agent repeats this process until either $x$ is completed or the maximum step limit $T$ is reached. Throughout the execution process, it is possible that some pairs of screenshots and instructions $(s_{t},x)$ may deviate significantly from the distribution of $\mathcal{D}_{\text{ID}}$. Such pairs are classified as OOD samples. When these OOD samples are encountered, the agent’s action predictions are prone to errors, which can lead to undesirable execution results.
The objective of OOD detection for GUI agents, therefore, is to identify whether the pair $(s_{t},x)$ at each time step $t$ deviates from the distribution of $\mathcal{D}_{\text{ID}}$. If OOD is detected, the agent should immediately terminate the execution of the instruction and alert the user; otherwise, it continues with the execution. To formally define the OOD detection mechanism, we introduce the following OOD detection function $f_{\text{OOD}}$, which operates on each pair $(s_{t},x)$:
$f_{\text{OOD}}(s_{t},x)=\begin{cases}1,&\text{if }(s_{t},x)\text{ is OOD},\\ 0,&\text{otherwise}.\end{cases}$ (1)
If $f_{\text{OOD}}(s_{t},x)$ returns a value of 1, indicating that the current state are OOD, the agent will terminate the process and alert the user. Otherwise, the agent will predict the appropriate action $a_{t}$ and proceed to execute it. The key to OOD detection for GUI agents is how to construct an effective $f_{\text{OOD}}(s_{t},x)$.
3.2 Pilot Study with Popular OOD Detection Methods
The existing popular OOD detection methods malinin2018predictive ; yang2024generalized can be broadly categorized into two main types: embedding-based methods lee2018simple ; sun2022out and uncertainty-based methods hendrycks2016baseline ; gal2016dropout . Based on existing research dong2024multiood ; lu2024recent in OOD detection for (M)LLMs, we used three embedding-based methods and two uncertainty-based methods for experimentation. Detailed descriptions of these methods can be found in Appendix E.
We evaluated the five OOD detection methods used as baselines across eight datasets spanning smartphone, computer, and web browser platforms. Following standard OOD detection method evaluation criteria luan2021out ; cui2022out , for each method, we plotted the ROC curve and reported the area under the receiver operating characteristic curve (AUROC) metz1978basic and false positive rate at 95% true positive rate (FPR95) for each device category dataset.
@@ 3.2 Pilot Study with Popular OOD Detection Methods
The pilot experiment results are shown in Table 1. The AUROC metric reflects the separability between OOD and ID datasets achieved by different methods in the GUI agent domain. Even the best baseline performances across different platforms only range from 76.14% to 89.77%, barely reaching the acceptable level traditionally expected for OOD detection. The FPR95 metric measures the false positive rate when the true positive rate reaches 95%. Pilot experiments show that even the best baselines result in an FPR95 of 31.75% to 48.92%, indicating that while maintaining 95% task success for the GUI agent, 31.75% to 48.92% of OOD tasks would still be mistakenly executed, posing significant risks.
Figure 2: TV score distribution on AndroidControl dataset and AITZ dataset.
@@
Figure 3: The output embeddings of the OS-Kairos dataset and AITZ dataset are visualized. Most of the sample points are confused together.
3.3 Why do these OOD detection methods perform suboptimally?
@@
For embedding-based methods, due to the relatively uniform reasoning patterns of GUI agents—TV score (as shown in Figure 3), which rely on differences in layer embeddings to infer reasoning path divergence, perform suboptimally in OOD detection for GUI agents. Other methods based on single-layer embeddings have achieved better performance; however, since GUI agents tend to lose some information related to input understanding during the reasoning process, there is still room to improve.
@@
As for uncertainty-based methods, the separability between the output spaces of ID and OOD samples in the GUI agent domain is extremely low (as shown in Figure 3). As a result, GUI agents struggle to estimate the uncertainty of their generated outputs, making uncertainty-based methods nearly ineffective at distinguishing between ID and OOD samples.
@@
These OOD detection methods rely on finding a decision boundary by identifying the Youden Index fluss2005estimation after obtaining some scoring metric, that is:
Youden $\text{Youden Index}=\arg\max_{t}\left(\text{TPR}(t)-\text{FPR}(t)\right),$ (2)
where $t$ is the threshold, and TPR and FPR represent the true positive rate and false positive rate at threshold $t$, respectively.
@@
However, popular GUI agents such as OS-Atlas wu2024atlas , UI-TARS qin2025ui , and AGUVIS xu2024aguvis have access to abundant and diverse training data. In the embedding space, the ID dataset representations for these GUI agents naturally form clusters according to their different data sources. Moreover, we observe that even for datasets originating from a single data source such as AITZ zhang2024android , as shown in Figure 4, there can still be noticeable clustering phenomena in the embedding space. We also observe that samples farther from the centroid tend to yield higher action success rates (see Appendix B.2). For such non-linearly separable data, the Youden Index approach becomes inadequate. Instead, models like GMM, which are capable of fitting distributions in a clustered manner, are better suited for our scenario.
4 GEM Method
Pilot experiments in Section 3 show that popular OOD detection methods perform suboptimally in the GUI agent domain. However, we observe that the input embeddings generated by GUI agents naturally lend themselves to modeling the ID representation of $\mathcal{D}_{\text{ID}}$. The algorithm description can be found in Appendix F.
Given a GUI agent $\mathcal{F}$ and an ID dataset $\mathcal{D}_{\text{ID}}=\{(s_{i},x_{i})\}_{i=1}^{k}$, we first obtain an encoder layer $l_{e}$ from $\mathcal{F}$. The encoder $l_{e}$ maps each input pair $(s_{i},x_{i})$ to an embedding vector $e_{i}\in\mathbb{R}^{n}$, formally:
$e_{i}=l_{e}(s_{i},x_{i}),\quad i=1,\ldots,k.$ (3)
Thus, we construct an ID embedding dataset $\mathcal{D}_{\text{embedding}}=\{e_{i}\}_{i=1}^{k}$. Each $e_{i}$ is a point in the $n$-dimensional embedding space.
To model this distribution, we first compute the centroid $\mu$ of $\mathcal{D}_{\text{embedding}}$: $\mu=\frac{1}{k}\sum_{i=1}^{k}e_{i}$. Next, we calculate the Euclidean distance between each embedding $e_{i}$ and the centroid $\mu$:
$d_{i}=\lVert e_{i}-\mu\rVert_{2},\quad i=1,\ldots,k,$ (4)
resulting in a distance dataset $\mathcal{D}_{\text{distance}}=\{d_{i}\}_{i=1}^{k}$.
The distribution of $\mathcal{D}_{\text{distance}}$ is typically nonlinear and may contain multiple modes. Therefore, instead of fitting a simple Gaussian or applying heuristic thresholds, we model it using a GMM. Specifically, we assume that the distances are generated from a mixture of $m$ univariate Gaussian components. The GMM models the probability density function as: $p(d)=\sum_{j=1}^{m}\pi_{j}\,\mathcal{N}(d\mid\mu_{j},\sigma_{j}^{2})$, $\pi_{j}$ is the mixing coefficient of the $j$-th component, satisfying $\sum_{j=1}^{m}\pi_{j}=1$ and $\pi_{j}\geq 0$, $\mathcal{N}(d\mid\mu_{j},\sigma_{j}^{2})$ denotes the density of a univariate Gaussian:
M-step update the parameters based on the estimated responsibilities:
$\pi_{j}^{\text{new}}\!\!=\frac{1}{k}\sum_{i=1}^{k}\gamma_{ij},\quad\mu_{j}^{\text{new}}\!\!=\frac{\sum_{i=1}^{k}\gamma_{ij}d_{i}}{\sum_{i=1}^{k}\gamma_{ij}},\quad\sigma_{j}^{2\,\text{new}}\!\!=\frac{\sum_{i=1}^{k}\gamma_{ij}(d_{i}-\mu_{j}^{\text{new}})^{2}}{\sum_{i=1}^{k}\gamma_{ij}}.$ (8)
$\text{BIC}(m)=-2\log\mathcal{L}_{m}+m\log k,$ (9)
The optimal number of components $m^{*}$ is thus selected as $m^{*}=\arg\min_{m}\text{BIC}(m)$.
Once the GMM is fitted with $m^{*}$ components, each Gaussian component $\mathcal{N}(\mu_{j},\sigma_{j}^{2})$ provides a center $\mu_{j}$ and a standard deviation $\sigma_{j}$. We define the ID boundary for each component as the interval $[\mu_{j}-n\sigma_{j},\mu_{j}+n\sigma_{j}]$. The details regarding the choice of $n$ can be found in Appendix G.
$f_{\text{OOD}}(s_{t},x)=\begin{cases}0,&\text{if }\exists j\text{ such that }d_{t}\in[\mu_{j}-n\sigma_{j},\mu_{j}+n\sigma_{j}],\\ 1,&\text{otherwise}.\end{cases}$ (10)
5 Experiments
5.1 Experiments Setup
Datasets. For the ID dataset, we use the AITZ zhang2024android dataset, which contains GUI agent data covering more than 70 Android app scenarios. For the OOD datasets, we select 8 GUI agent datasets spanning three platforms: smartphone, computer, and web browser. The smartphone platform includes AndroidControl rawles2024androidworld , OS-Kairos cheng2025kairos , Meta-GUI sun2022meta , and ScreenSpot-Mobile li2025ScreenSpot ,while the computer platform includes Omniact-Desktop kapoor2024omniact and ScreenSpot-Desktop, and the web browser platform includes Omniact-Web and ScreenSpot-Web. The four smartphone datasets simulate extrapolation-OOD scenarios, whereas the computer and web browser datasets simulate internalization-OOD scenarios. Detailed description of the datasets is provided in Appendix C.
Table 2: Comparison of OOD Detection Results Across Different Datasets. For each metric, the best result among baseline methods is underlined.
Method Metric AndroidControl OS-Kairos
Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
@@ 5.1 Experiments Setup
Last layer embedding 70.08 69.02 96.77 80.57 32.66 17.45 91.10 29.29
Best layer embedding 78.20 77.02 94.07 84.69 74.35 33.29 67.33 44.56
@@
[…]
Method Metric Meta-GUI ScreenSpot-Mobile
Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
@@
Last layer embedding 59.38 39.47 75.72 51.89 30.71 12.05 98.61 21.47
Best layer embedding 76.59 55.50 96.31 70.42 62.72 19.60 92.83 32.36
@@
[…]
Method Metric Omniact-Desktop ScreenSpot-Desktop
Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
@@
Last layer embedding 60.05 35.80 83.84 50.17 46.54 10.05 89.22 18.06
Best layer embedding 91.15 83.73 78.34 80.94 43.46 10.18 96.71 18.43
@@
[…]
Method Metric Omniact-Web ScreenSpot-Web
Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
@@
Last layer embedding 61.82 18.52 81.89 30.20 47.77 12.64 87.61 22.09
Best layer embedding 86.03 40.34 80.38 53.72 74.77 24.77 97.48 39.50

### v3

@@ Abstract
Graphical user interface (GUI) agents have recently emerged as an intriguing paradigm for human-computer interaction, capable of automatically executing user instructions to operate intelligent terminal devices. However, when encountering out-of-distribution (OOD) instructions that violate environmental constraints or exceed the current capabilities of agents, GUI agents may suffer task breakdowns or even pose security threats. Therefore, effective OOD detection for GUI agents is essential. Traditional OOD detection methods perform suboptimally in this domain due to the complex embedding space and evolving GUI environments. In this work, we observe that the in-distribution input semantic space of GUI agents exhibits a clustering pattern with respect to the distance from the centroid. Based on the finding, we propose GEM, a novel method based on fitting a Gaussian mixture model over input embedding distances extracted from the GUI agent that reflect its capability boundary. Evaluated on eight datasets spanning smartphones, computers, and web browsers, our method achieves an average accuracy improvement of 23.70% over the best-performing baseline while only increasing training time by 4.9% and testing time by 6.5%. We also experimentally demonstrate that GEM can improve the step-wise success rate by 9.40% by requesting assistance from the cloud model when encountering OOD samples. Analysis verifies the generalization ability of our method through experiments on nine different backbones. The codes are available at https://github.com/Wuzheng02/GEM-OODforGUIagents.
@@
Introduction
Recently, graphical user interface (GUI) agents (Zhang et al. 2024a; Liu et al. 2025b; Shi et al. 2025) have emerged as an intriguing paradigm to human-computer interaction, capable of autonomously executing user instructions and performing human-like control on intelligent terminal devices such as smartphones, computers, and web browsers. The common approach to building GUI agents involves post-training multimodal large language models (MLLMs) using high-quality trajectory data to enhance key task capabilities such as perception (Li et al. 2023; Wang et al. 2023b), reasoning (Yao et al. 2023; Wei et al. 2022), and reflection (Liu et al. 2025c; Hu et al. 2025). Despite notable advances in instruction following, GUI agents remain vulnerable to out-of-distribution (OOD) risks—executing instructions that violate environmental constraints (e.g., non-existent functions or applications) or exceed the agent’s current capabilities. These failures can lead to task breakdowns or even pose security threats. In real-world applications, OOD risks for GUI agents come with two primary forms: (i) Internalization-OOD, where the agent operates in domain-specific environments (e.g., a particular type of smartphones or vehicle cabins) but incorrectly assumes the presence of unsupported capabilities; (ii) Extrapolation-OOD, where the agent encounters instructions tied to dynamic or evolving environments (e.g., new third-party applications).
@@ Introduction
Popular OOD detection methods can be broadly categorized (Malinin and Gales 2018) into two types: embedding-based approaches (Lee et al. 2018) and model uncertainty-based approaches (Hendrycks and Gimpel 2016). These methods have demonstrated effectiveness in traditional (M)LLM tasks such as summarization (Ren et al. 2022), visual question answering (Kervadec et al. 2021), mathematical reasoning (Wang et al. 2024b), and text generation (Wu et al. 2022). Directly applying these existing OOD detection techniques to the GUI agent domain results in suboptimal performance. Notably, even with the classification threshold calibrated to correctly identify 95% of in-distribution (ID) tasks, the best-performing method still misclassifies between 31.75% and 48.92% of OOD tasks as ID.
@@
(i) Complex Embedding Space: The inputs to GUI agents are inherently complex (Ma, Zhang, and Zhao 2024). Concretely, GUI screens typically contain densely populated UI components (Zhang et al. 2024a), leading to higher information density than traditional MLLM tasks. Besides, the diversity of user instructions further complicates the inference of user intent. This complexity substantially increases the difficulty of effective OOD detection for GUI agents.
(ii) Evolving Environments: GUI environments frequently change due to system upgrades or the installation of new third-party applications. As a result, GUI agents must continually adapt their capabilities (Wang et al. 2025; Liu et al. 2025a). This evolving nature complicates OOD detection by demanding both accurate capability assessment and temporal adaptability. Consequently, reliance on static external classifiers becomes increasingly impractical.
To address the challenges above, we propose GEM, a novel method based on fitting a Gaussian mixture model (GMM) (Reynolds et al. 2009) over input embedding distances extracted from the GUI Agent that reflect its capability boundary. Specifically, we first extract the input embeddings from the encoder layer of the GUI agent on its training data, and fit them into a high-dimensional hypersphere. We then compute the L2-norm distances of all embeddings relative to the centroid of the hypersphere. Under bayesian information criterion (BIC) (Neath and Cavanaugh 2012) supervision, we fit a GMM to these distances and define the OOD detection boundary as a configurable number of standard deviations away from each GMM cluster center. Experiments across eight GUI agent datasets spanning smartphones, computers, and web browsers platforms show that our method consistently outperforms all traditional OOD detection methods. And we provide a possible method to use GEM to improve the end-to-end results for GUI agents.
@@
(ii) We propose a novel GMM-based approach that leverages input embedding distances from MLLMs for OOD detection. Our method achieves an average accuracy improvement of 23.70% over the best performing baseline across eight datasets spanning three platform types while only increasing training time by 4.9% and testing time by 6.5%.
(iii) We offer some new insights into this emerging research area and provide a possible method to use GEM to improve the end-to-end results for GUI agents.
@@
Related Work
MLLM-based GUI Agents
With the advancement of MLLMs (Liu et al. 2023; Achiam et al. 2023), numerous GUI agent foundation models (Hong et al. 2024; Wu et al. 2024; Qin et al. 2025; Xu et al. 2024) have emerged, capable of executing user instructions across smartphones, computers, and web browsers. For the technical framework, prompt-based approaches (Wang et al. 2025; Jiang et al. 2025) have been developed, leveraging closed-source models to construct agent systems that fulfill user instructions. Additionally, researchers have also explored pre-training (Ma, Zhang, and Zhao 2024; Wu et al. 2025b), supervised fine-tuning (SFT) (Zhang and Zhang 2024; Wu et al. 2025a), and reinforcement learning (RL) (Zhou et al. 2024; Liu et al. 2024; Wang et al. 2024a; Xia and Luo 2025) techniques to further enhance the ability of GUI agents to complete user instructions. Regardless of their advancements, GUI agents inevitably face capability limitations. In complex and dynamic real-world environments (Cheng et al. 2025a; Guo et al. 2025), GUI agents are prone to encountering OOD situations.
OOD Detection in MLLMs
For OOD detection in MLLMs, it is necessary to jointly consider both visual and textual modalities, in contrast to traditional computer vision OOD detection, which only requires modeling the visual modality (Liu et al. 2020; Ren et al. 2019), or LLMs OOD detection, which focuses on textual modality (Wang et al. 2024b; Ren et al. 2022). In the context of MLLMs, some researchers have proposed using maximum concept matching (Ming et al. 2022) to characterize OOD uncertainty, while others have approached the problem by training models on ID datasets and classifying whether an input image belongs to an unknown category (Wang et al. 2023a). However, OOD detection for MLLMs remains a challenging research area (Dong et al. 2024; Lu et al. 2024). Furthermore, for MLLM-based GUI agents, task scenarios are significantly more complex and dynamic (Shi et al. 2025) compared to traditional MLLM tasks, making OOD detection even more difficult.
Investigating the Challenge of OOD Dection for GUI Agents
In this section, we conduct pilot experiments to evaluate the effectiveness of popular OOD detection methods for GUI agents and understand the challenges. We will provide the problem formulation and present the experimental settings followed by the key results and analysis of the pilot study.
Problem Formulation
A GUI agent $\mathcal{F}$ is initially trained on an ID dataset $\mathcal{D}_{\text{ID}}=\{(s_{i},x_{i})\}_{i=1}^{k}$, consisting of $k$ pairs of screenshots $s_{i}$ and user instructions $x_{i}$. $\mathcal{F}$ learns knowledge of operating systems through methods such as SFT and RL on $\mathcal{D}_{\text{ID}}$.
After deployment on the device, $\mathcal{F}$ receives an instruction $x$ and captures the current device screenshot $s_{t}$. $\mathcal{F}$ then constructs a prompt that combines $s_{t}$ and $x$, which is subsequently used to predict an action $a_{t}$. Formally, at each time step $t$ ($0\leq t\leq T$), the process can be expressed as: $a_{t}=\mathcal{F}(s_{t},x)$ where $s_{t}$ is the screenshot at time step $t$, $x$ is the instruction, and $a_{t}$ is the action predicted by the agent.
Then $a_{t}$ is executed, leading to a new screenshot $s_{t+1}$. $\mathcal{F}$ evaluates whether the instruction $x$ has been completed. If $x$ is not completed, the agent repeats this process until either $x$ is completed or the maximum step limit $T$ is reached. Throughout the execution process, it is possible that some pairs of screenshots and instructions $(s_{t},x)$ may deviate significantly from the distribution of $\mathcal{D}_{\text{ID}}$. Such pairs are classified as OOD samples. When these OOD samples are encountered, the agent’s action predictions are prone to errors, which can lead to undesirable execution results.
The objective of OOD detection for GUI agents, therefore, is to identify whether the pair $(s_{t},x)$ at each time step $t$ deviates from the distribution of $\mathcal{D}_{\text{ID}}$. If OOD is detected, the agent should immediately terminate the execution of the instruction and alert the user; otherwise, it continues with the execution. To formally define the OOD detection mechanism, we introduce the following OOD detection function $f_{\text{OOD}}$, which operates on each pair $(s_{t},x)$.
If $(s_{t},x)\text{ is OOD}$, $f_{\text{OOD}}(s_{t},x)$ returns a value of 1, indicating that the current state are OOD. Otherwise, the agent will predict the appropriate action $a_{t}$ and proceed to execute it. The key to OOD detection for GUI agents is how to construct an effective $f_{\text{OOD}}(s_{t},x)$.
Pilot Study with Popular OOD Detection Methods
The existing popular OOD detection methods (Malinin and Gales 2018; Yang et al. 2024) can be broadly categorized into two main types: embedding-based methods (Lee et al. 2018) and uncertainty-based methods (Gal and Ghahramani 2016). Based on existing research (Dong et al. 2024; Lu et al. 2024) in OOD detection for (M)LLMs, we used three embedding-based methods and two uncertainty-based methods for experimentation.
We evaluated the five OOD detection methods used as baselines across eight datasets spanning three platforms. Following standard OOD detection method evaluation criteria (Luan et al. 2021; Cui and Wang 2022), we plotted the ROC curve and reported the area under the receiver operating characteristic curve (AUROC) (Metz 1978) and false positive rate at 95% true positive rate (FPR95).
@@
The pilot experiment results are shown in Table 1. The AUROC metric reflects the separability between OOD and ID datasets achieved by different methods. Even the best baseline performances only range from 76.14% to 89.77%, barely reaching the acceptable level expected for OOD detection. The FPR95 metric measures the false positive rate when the true positive rate reaches 95%. Pilot experiments show that even the best baselines result in an FPR95 of 31.75% to 48.92%, indicating that while maintaining 95% task success for the GUI agent, 31.75% to 48.92% of OOD tasks would be mistakenly executed, posing significant risks.
@@
Figure 2: The output embeddings of the OS-Kairos and AITZ are visualized. Most of the samples are confused.
Why do these OOD detection methods perform suboptimally?
@@
For embedding-based methods, due to the relatively uniform reasoning patterns of GUI agents—TV score, which rely on differences in layer embeddings to infer reasoning path divergence, perform suboptimally. Other methods based on single-layer embeddings have achieved better performance; however, since GUI agents tend to lose some information related to input understanding during the reasoning process, there is still room to improve.
@@
For uncertainty-based methods, the separability between the output spaces of ID and OOD samples in the GUI agent domain is extremely low (as shown in Figure 2). As a result, GUI agents struggle to estimate the uncertainty of their outputs, making uncertainty-based methods nearly ineffective to distinguish between ID and OOD samples.
@@
These OOD detection methods rely on finding a decision boundary by identifying the Youden Index (Fluss, Faraggi, and Reiser 2005) after obtaining some scoring metric, that is $\text{Youden Index}=\arg\max_{t}\left(\text{TPR}(t)-\text{FPR}(t)\right)$, where $t$ is the threshold, and TPR and FPR represent the true positive rate and false positive rate.
@@
However, popular GUI agents (Wu et al. 2024; Qin et al. 2025; Xu et al. 2024) have access to diverse training data. In the embedding space, the ID dataset representations for these GUI agents naturally form clusters according to their different data sources. Moreover, we observe that even for datasets originating from a single data source such as AITZ (Zhang et al. 2024b), as shown in Figure 3, there can still be noticeable clustering phenomena. We also observe that samples farther from the centroid tend to yield higher success rates (see Appendix). For such non-linearly separable data, the Youden Index approach becomes inadequate.
GEM Method
AndroidControl OS-Kairos
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Last layer embed. 70.08 69.02 96.77 80.57 32.66 17.45 91.10 29.29
Best layer embed. 78.20 77.02 94.07 84.69 74.35 33.29 67.33 44.56
@@
Meta-GUI ScreenSpot-Mobile
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Last layer embed. 59.38 39.47 75.72 51.89 30.71 12.05 98.61 21.47
Best layer embed. 76.59 55.50 96.31 70.42 62.72 19.60 92.83 32.36
@@
Omniact-Desktop ScreenSpot-Desktop
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Last layer embed. 60.05 35.80 83.84 50.17 46.54 10.05 89.22 18.06
Best layer embed. 91.15 83.73 78.34 80.94 43.46 10.18 96.71 18.43
@@
Omniact-Web ScreenSpot-Web
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Last layer embed. 61.82 18.52 81.89 30.20 47.77 12.64 87.61 22.09
Best layer embed. 86.03 40.34 80.38 53.72 74.77 24.77 97.48 39.50